What is the compressed air pressure in a 80 liter capacity cylinder if the temperature is 15 degrees and the mass is 4 kilograms?
1
Expert's answer
2021-04-05T11:08:25-0400

The pressure can be expressed from the following equation:


"pV=\frac{m}{\Mu}RT,\\\space\\
p=\frac{m}{V\Mu}RT,\\\space\\
p=\frac{4}{(80·10^{-3})(29·10^{-3})}8.314(15+273.15)=\\\space\\=41.31·10^6\text{ Pa}."